OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 12 and the last term is -3,359,232? -3,599,175 -3,359,230 -3,119,285 -2,879,340

OpenStudy (anonymous):

\[ 12 a^7=-3359232 \\ a=- 6 \]

OpenStudy (anonymous):

The terms are {12, -72, 432, -2592, 15552, -93312, 559872, -3359232}

OpenStudy (anonymous):

Their sum is -2879340

OpenStudy (anonymous):

You can also apply the formula \[ \frac{12 \left(1-a^8\right)}{1-a}=\frac{12 \left(1-(-6)^8\right)}{1+6}=-2879340 \]